eBook Exercise 9.13 (Algorithmic) Consider the following hypothesis test: но: 50 Ha: μ > 50 A sample of 55 is used and the population standard deviation is 7. Use the critical value approach to state your conclusion for each of the following sample results. Use a-.05 a. With x = 52.5, what is the value of the test statistic (to 2 decimals)? Can it be concluded that the population mean is greater than 50? Select your answer b. With x-51, what is the value of the test statistic (to 2 decimals)? Can it be concluded that the population mean is greater than 50? - Select your answer C. With x = 51.8, what is the value of the test statistic (to 2 decimals)? Can it be concluded that the population mean is greater than 50? Select your answer-
Answer #1

SolutionA:

z=xbar-mu/sigma/sqrt(n)

z=(52.5-50)/(7/sqrt(55))

z=2.65

value of the test statistic:z=2.65

since critical z=1.96

test statistic>critical z value

yes,it be concluded that population mean is greater than 50.

ANSWERS:

value of the test statistic:z=2.65

YES

Solutionb:

z=(51-50)/(7/sqrt(55))

z=1.06

value of the test statistic:z=1.06

since critical z=1.96

test statistic<critical z value

no,it cannot be concluded that population mean is greater than 50.

ANSWER

value of the test statistic:z=1.06

NO

Soutionc:

z=(51.8-50)/(7/sqrt(55))

z=1.91

test statistic<critical z value

no,it cannot be concluded that population mean is greater than 50.

ANSWER

value of the test statistic:z=1.91

NO